Air contains23%oxygenand77%nitrogen by weight the percentage  of oxygen in volume is???

Grade:

1 Answers

Arun
5 years ago
Assume that the calculations are for 100 g of air:-

Mass of oxygen (O₂) in 100 g of air = 23 g

Mass of nitrogen (N₂) in 100 g of air = 77 g

Molar mass of Oxygen = 32 g

1 mole of oxygen or 32 g of Oxygen occupies 22.4 L volume.
 
Therefore, volume of 23 g of oxygen will occupy 22.4/32 x 23 = 16.1 L

Similarly, 28 g of Nitrogen occupies 22.4 L volume.
 
Therefore, 77 g of Nitrogen will occupy 22.4/28 x 77 = 61.6 L

Total volume of air = 16.1 + 61.6 = 77.7 L

% by volume of Oxygen = 16.1/77.7 x 100 = 20.7 %